OpenStudy (anonymous):

Find the equation of a line with the given points. (Put your answers in the form y = mx + b.) (2, 2) (3, 3)

OpenStudy (anonymous):

find the slope and then use it and one of the points to find the equation of the line. Post your answer and we will let you know if you are correct

OpenStudy (anonymous):

y=x

OpenStudy (anonymous):

its obvious :|

OpenStudy (anonymous):

\[(y-2)/(x-2) = (3-2)/(3-2)\] \[y = x\]

OpenStudy (anonymous):

you find the slope first by finding rise over run. Slope equals m. m= (y2-y1)/(x2-x1) m= .1/1 or 1 (for every unit to the right the line moves one unit up) Now plug in m and your points into y=mx+b. I will use (2,2) 2=m(2) + b 2=1(2)+b 2=2+b (subtract 2 from each side) 0=b

OpenStudy (anonymous):

so line equals y=x